MsBrains (ms-brains):

@kayders1997 Explaining this to me

MsBrains (ms-brains):

I think it's C.-48

OpenStudy (kayders1997):

So you still have xk=y so what would you plug into those values for the first part?

OpenStudy (mathmale):

It'd be so much more helpful if you'd explain your reasoning instead of saying, "I think it's c." kayders1997 is correct in bringing up the formula for direct variation, y=kx. Thank you.

MsBrains (ms-brains):

Actually, I'm not sure why I think it's C. :0 We would plug in -4, and 12, right?

OpenStudy (kayders1997):

No -4K=? what's the other value that corresponds to that part

OpenStudy (mathmale):

Ms-Brains: Think back to what we've done before. Apply what you've learned earlier. How did we find the value of k earlier? Do the same thing here. What is k? Then, what is the whole "direct variation" equation?

OpenStudy (kayders1997):

24 or 12?

MsBrains (ms-brains):

k=12?

OpenStudy (kayders1997):

Because it says f(x)=24 when x is-4 were going to use 24 the others number is for the next part

OpenStudy (kayders1997):

Y=24 x=4 you have to put that into the equation xk=y

MsBrains (ms-brains):

(4)12=24 <---like that?

OpenStudy (kayders1997):

No we don't know k because they don't give it to us in and -4 my bad but otherwise it looks good

OpenStudy (kayders1997):

We're trying to find k

MsBrains (ms-brains):

-24 = 24 k=-2

OpenStudy (kayders1997):

No x=-4 y=24 put it into equation xk=24 -4(k)=24 do you understand where these numbers come from?

OpenStudy (mathmale):

We would plug in -4, and 12, right? No. Instead, plug in x=-4 and y=24 in y=kx. Find k.

OpenStudy (mathmale):

y = k x Let x=-4 and y=24. Find k. Now write y = k x, this time including the value for k that you've just found.

MsBrains (ms-brains):

24=(x)-4 <- like this?

OpenStudy (kayders1997):

X is -4 so kinda close

OpenStudy (kayders1997):

The 24 is right though

OpenStudy (mathmale):

You appear to be subtracting. There is no subtraction in this problem. Fix your equation, please.

OpenStudy (kayders1997):

Rememeber were trying to find k

OpenStudy (mathmale):

"kinda close" doesn't work here. We need to be exact.

OpenStudy (kayders1997):

You gotta have some positivity not everyone gets things in a fast rate

OpenStudy (kayders1997):

Making mistakes is the best way to learn what you are doing wrong

MsBrains (ms-brains):

if not subtraction, what are we using? multiplication?

OpenStudy (kayders1997):

Yes

OpenStudy (kayders1997):

Xk that's multiplication just like 3(2) mutliplication

OpenStudy (mathmale):

y = k x involves only multiplication. No subtraction. You could fix your equation by inserting parentheses in the proper places.

OpenStudy (mathmale):

24=(x)-4 <- like this?

MsBrains (ms-brains):

24=(k)*4 Okay, so we have the equation. Now what?

OpenStudy (kayders1997):

-4 for x though now solve for k by?

OpenStudy (mathmale):

y = k x Here, x=-4 and y=+24. Find k. k will be negative.

OpenStudy (kayders1997):

How would you get rid of -4K to just get k alone?

OpenStudy (kayders1997):

Or what times -4 is 24 another way to think of it

OpenStudy (mathmale):

Model: y = k x Values of x and y: x=(-4) and y=+24 Substitute these values into y = k x now. Then we'll solve the resulting equation for k.

MsBrains (ms-brains):

24=(-6)*-4 24=24 So -6 is the correct answer?

OpenStudy (mathmale):

Yes. But would you follow the plan I suggested a minute ago? +24 = K ( ? )

OpenStudy (kayders1997):

Yeah that the first part now we have to plug in the 12 for x so we know k is -6 so plug in values for k and x Kx=y and what does y end up being?

MsBrains (ms-brains):

-6 isn't the final answer?

OpenStudy (kayders1997):

No because it's asking what is y when x=12 for the second part by k is -6

OpenStudy (kayders1997):

But

OpenStudy (mathmale):

Indeed it is not; you have to find the proportionality constant, k, first. k = -6. Now write out the direct variation formula including this k = -6.

OpenStudy (kayders1997):

So all you have to do it plug in the numbers you know k=-6 x=12 Xk=y what is y?

OpenStudy (kayders1997):

And that will be your answer :)

MsBrains (ms-brains):

24(12)-6 write the final formula like this..

OpenStudy (kayders1997):

What do you plug in for x and k?

OpenStudy (mathmale):

In summary: 1. Use the given info to find the proportionality constant, k. It's -6. 2. Write out the complete equation for direct variation that applies here. Because k = -6, it is y = -6x 3. Now, in a particular case, where another x is given, find the resulting y value.

OpenStudy (kayders1997):

24 is not in this

OpenStudy (kayders1997):

12(-6) X(k)

OpenStudy (mathmale):

Thank you for showing your work: 24(12)-6 write the final formula like this.. Actually, the final formula (not the final answer) is y = -6x. Let x=12 (a particular x-value). Subst. 12 for x in the above equation and calculate y.

OpenStudy (kayders1997):

X=12 k=-6

MsBrains (ms-brains):

Alright. so the final answer is -72. Thank you both for teaching me how to do it. ^-^ You guys rock!

OpenStudy (mathmale):

You're welcome. Yes, when x = 12, y = -72. Very good.